\begin{thebibliography}{99}

\bibitem{b2}
Bernstein D. J. {\it Removing redundancy in high-precision Newton
iteration.} Manuscript, 2004. {\sf
http://cr.yp.to/papers.html\#fastnewton.}

\bibitem{br}
Brent R. {\it Multiple-precision zero-finding methods and the
complexity of elementary function evaluation.} Analytic
computational complexity. Academic Press, NY, 1975, 151--176.

\bibitem{gg}
von zur Gathen J., Gerhard J. {\it Modern computer algebra.}
Cambridge University Press, Cambridge, 1999.

\bibitem{h}
Harvey D. {\it Faster exponentials of power series.} Preprint.
2009. {\sf http://arxiv.org/abs/0911.3110.}

\bibitem{vdh}
van der Hoeven J. {\it Newton's method and FFT trading.} Tech.
report 2006--17. Univ. Paris-Sud, Orsay, France, 2006.

\bibitem{vdh2}
van der Hoeven J. {\it Newton's method and FFT trading.} J. Symb.
Comput. 2010. {\bf 45}(8), 857--878.

\bibitem{km}
Karp A. H., Markstein P. {\it High-precision division and square
root.} ACM Trans. Math. Softw. 1997. {\bf 23}(4), 561--589.

\bibitem{iss}
Sergeev I. S. {\it Fast algorithms for elementary operations on
power series.} Proc. IX Intern. Seminar ``Discrete math. and its
applications'' $($Moscow, June $2007$). Moscow, MSU Mech. Math.
Faculty, 2007, 123--126 (in Russian).

\bibitem{is9}
Sergeev I. S. {\it Regularization of some estimates of complexity
of the polynomial multiplication.} Proc. VII Sci. Workshop on
Discrete math. and its applications (Moscow, May 2009). Part II.
Moscow, Keldysh Inst. of Applied Math., 2009, 26--32 (in Russian).

\bibitem{is10}
Sergeev I. S. {\it Fast algorithms for elementary operations on
complex power series.} Discrete Math. and Appl. 2010. {\bf 20}(1),
25--60.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{latexsym,amssymb,amsmath}

\begin{document}
\title{A note on the fast power series' exponential}
\date{}
\author{I. S. Sergeev}

\maketitle
\begin{abstract}
It is shown that the exponential of a complex power series up to
order $n$ can be implemented via $(23/12+o(1))M(n)$ binary
arithmetic operations over $\mathbb C$, where $M(n)$ stands for
the (smoothed) complexity of multiplication of polynomials of
degree $<n$ in FFT-model. Yet, it is shown how to raise a power
series to a constant power with the complexity $(27/8+o(1))M(n)$.
\end{abstract}

\section{Introduction}

It is very well known that the exponential of a power series (as
well as some other elementary operations) has the same order of
complexity as multiplication, see e.g.~\cite{br,gg}. (When we
speak about complexity we consider a computational model of
circuits or straight-line programs over arithmetic basis $\{ \pm,
* \} \cup \{ ax \mid a \in \mathbb C \}$, see e.g.~\cite{gg}. The field should not be
necessary complex, it might be real or any algebraically closed,
or any other which supports appropriate FFT.)

Previous papers set up a convention to estimate complexity of the
basic power series' operations (including exponential) in the
number of multiplications of the same size. The very last papers
(since 2000) use the special FFT-way multiplications.

The complexity function $M(n)$ of FFT-way multiplication can be
introduced as follows. Let $F^{\circ}(n)$ be the complexity of FFT
of order $n$ over $\mathbb C$. Let $F(n)$ be its smoothed version,
that is, $F(n) = \min_{m \ge n} F^{\circ}(m)$. For simplicity we
also assume $F(n)=\Omega(n\log n)$ (in fact, $F(n)=\omega(n)$ is
sufficient) and $kF(n)=(1+o(1))F(kn)$ for $1 \le k \le \log
n$.\footnote{These assumptions are for more convenient way of
writing final complexity bounds only, it does not affect any other
aspect of the proof. Notation $g=\omega(f)$ means $f=o(g)$.} Then,
let $M(n) = 6F(n)$. In any case $M(n)$ serves as an upper
asymptotic estimate of the complexity of multiplication of
polynomials of degree $<n$.

Write an upper estimate of the complexity of computing exponent of
a power series in $\mathbb C[[x]]$ modulo $x^n$ in the form
$(A+o(1))M(n)$. Let us list some previously known results:
$A=17/6$~\cite{b2}, $A=14/3$~\cite{vdh,vdh2},
$A=13/6$~\cite{iss,is10} and~\cite{h}. Next, we will show that
$A=23/12$ is also valid. The method is a straightforward
combination of methods~\cite{iss,is10} and~\cite{h}. To be more
precise, scheme of computation follows~\cite{h} and technique is
inherited from~\cite{iss,is10}.

We will also show that one can compute a constant power of a power
series in $\mathbb C[[x]]$ modulo $x^n$ with the complexity
$(B+o(1))M(n)$, where $B=27/8$. It slightly improves the
previously best known factor $B=41/12$~\cite{iss,is10}.

Some notation. Let $f \in \mathbb C[[x]]$, then $f_{..n}$ denotes
$f \bmod x^n$ and $\lfloor f/x^n \rfloor$ stands for
$(f-f_{..n})/x^n$. If $f = \sum_{i \ge 0} f_i x^i$, then $\Delta
f$, ${\bf J} f$, $\ln f$ (when $f_0=1$) and $e^f$ (when $f_0=0$)
denotes formal derivative, formal integral, formal logarithm and
formal exponent respectively:
$$ \Delta f = \sum_{i \ge 1} if_i x^{i-1}, \quad {\bf J} f = \sum_{i \ge 1} \frac{f_{i-1}}i x^{i},
\quad  \ln f = - \sum_{i \ge 1} \frac{(1-f)^i}{i},  \quad e^f =
\sum_{i \ge 0} \frac{f^i}{i!}. $$

\section{Exponent}

Consider a problem of computing exponent of a power series $h$,
$h_{..1}=0$. Denote $f=e^h$, $r=1/f$. Recall that $\Delta h=\Delta
f/f$.

The next iterative formula~\cite{vdh} (derived as a solution of an
equation $\Delta f = g f$ with $g=\Delta h$ in this case) is valid
for $m\ge n$:
\begin{equation}
f_{..m+n} = f_{..m} + f_{..n} \, {\bf J} \left( x^{m-1}r_{..n}
\left\lfloor \Delta (h_{..m+n}) f_{..m} / x^{m-1} \right\rfloor
\right) \mod x^{m+n}.
\end{equation}

Let $E(n)$ and $I(n)$ denote the complexity of computation $e^f$
and $1/f$ modulo $x^n$ respectively. Then we can use~(1) to
compute $f_{..m}$ with the complexity
\begin{equation}
E(n) + I(n) + (13+o(1))F(m) \sim (13/6+o(1))M(m)
\end{equation}
for appropriately chosen parameters $m$ and $n$, e.g. $n=o(m)$ and
$m=O(n\sqrt{\log n})$. (That is one of the ways to obtain factor
$13/6$ in the complexity estimate for exponent~\cite{is10}.)

To achieve the complexity bound~(2) split series into blocks of
appropriate size $k$, e.g. $k \in o(n) \cap \Omega(n/\sqrt{\log
n})$. Then use double DFT of order $(2k,k)$ to perform block
multiplications in~(1).

{\it Double DFT} of order $(l,k)$ as a map from $\mathbb C[x]$ to
$\mathbb C^{l+k}$ is defined so that its first $l$ components are
the components of DFT of order $l$, another $k$ components are the
components of composition of the variable substitution $x \to
\zeta x$ and DFT of order $k$, where $\zeta$~--- an appropriate
complex number. Multiple DFT can be defined in similar
way~\cite{is9}. Multiple DFTs are useful to perform
multiplications of different sizes on the overlapping sets of
inputs. Double DFT of order $(l,k)$ or its inverse costs as much
as DFT of order $l$ and DFT of order $k$ plus $O(l+k)$ extra
operations, see~\cite{is10} for details.

In the case $l=2k$ (as we have) one can use an ordinary DFT of
order $3k$ decomposed into outer DFTs of order $3$ and inner DFTs
of order $k$ instead.

The main term of the complexity estimate~(2) is contributed by:
\linebreak $3(m/k)F(k)$ (which we assume to be approximately
$3F(m)$) operations to compute DFTs of blocks of $f$, the same
number of operations to compute DFTs of blocks of $\Delta h$, the
same number of operations to restore blocks of the triple product
under the integral, $2F(m)$ operations to compute $2k$-order parts
of DFTs of blocks of ${\bf J}(\ldots)$ and the same number of
operations to restore the product $f_{..n} \, {\bf J} (\ldots)$.
Other steps (precomputation of $f_{..n}$ and $r_{..n}$, additions,
implementation of $\Delta$ and $\bf J$ operators, calculations in
the DFT-image spaces) contribute $o(m\log m)$ in total complexity.

To provide a hint for verification we consider a subproblem of the
triple product computation (this step seems to be less evident in
the algorithm above) in Appendix. Other details (if necessary) see
in~\cite{is10}.

Next we turn to introduce an improved version of the algorithm
with the use of idea due to D.~Harvey~\cite{h}.

Suppose we are to compute $f_{..2m}$. Firstly we compute $f_{..m}$
acting as mentioned above. By the way we also have DFTs of blocks
of $f_{..m-n}$ and $\Delta (h_{..m})$ been computed. At the second
stage we use formula~\cite{br}
\begin{equation}
f_{..2m} = f_{..m} + f_{..m}(h - \ln f_{..m})_{..2m} \mod x^{2m}
\end{equation}
derived by the discrete Newton---Raphson method as the solution of
an equation $\varphi[f]=h$ with $\varphi[x]= \ln x$ in our case.
This stage generally consists of the two essential parts:
computation of $\ln f_{..m}$ up to order $2m$ and final
multiplication $f_{..m}$ by $h-\ln f_{..m}$.

Denote $s=\Delta(f_{..m})/f_{..m}$. We compute $(\ln
f_{..m})_{..2m}$ as ${\bf J} s_{..2m-1}$ via the
iteration~\cite{km}
\begin{equation}
 s_{..m'+n-1} = s_{..m'-1} - x^{m'-1}r_{..n}\left\lfloor s_{..m'-1}f_{..m} / x^{m'-1} \right\rfloor \mod x^{m'+n-1},
\end{equation}
where $m'\ge m$. We start from $s_{..m-1} = \Delta (h_{..m})$.

To perform the calculations by~(4) we use $(6+o(1))F(m)$ extra
operations: half of them to compute DFTs of remaining blocks of
$s$, another half to compute blocks of the triple product (see
Appendix for some details).

To complete the computation of $f_{..2m}$ we use another
$(4+o(1))F(m)$ operations: half of them to compute $2k$-order DFTs
of the order of blocks of $h-\ln f_{..m}$ and roughly the same
number to restore blocks of the product in~(3). Recall that
$2k$-order DFTs of almost all blocks of $f_{..m}$ are also
computed at the first stage of the algorithm since we use double
DFTs (this is the only place we gain a benefit from exploiting
double DFT).

To summarize, we can compute the exponent up to order $2m$ with
the complexity $(23+o(1))F(m)$.

\section{Exponentiation}

Consider a problem of raising of a power series $h$, $h_{..1}=1$
to a power $C \in \mathbb C$. Denote $f=h^C$, $r=1/f$, $\rho =
1/h$, $s=C\Delta h/h$.

The way of computing a power is just similar with that of
computing an exponent. We will give only a sketch.

To compute the first half of the required coefficients of $f$ we
use the formula
\begin{equation}
f_{..m+n} = f_{..m} + f_{..n} \, {\bf J} \left( x^{m-1} r_{..n}
\left\lfloor s_{..m+n-1}f_{..m} / x^{m-1} \right\rfloor \right)
\mod x^{m+n}
\end{equation}
derived as a solution of the equation $\Delta f = s f$. Next, we
switch to the formula
\begin{equation}
f_{..2m} = f_{..m} + f_{..m}({\bf J} s - \ln f_{..m})_{..2m} \mod
x^{2m}
\end{equation}
derived as a solution of the equation $\ln f = C\ln h$.

To solve a subproblem of computing $s$ we use the
iteration~\cite{km}
\begin{equation}
 s_{..m+n-1} = s_{..m-1} + \rho_{..n}(\Delta(h_{..m+n}) - s_{..m-1}h_{..m+n}) \bmod x^{m+n}.
\end{equation}

As above all series are divided into blocks of size $k$, except
that the first half of the series $\Delta h$ is divided into
blocks of size $2k$. We also use double DFTs of order $(2k,k)$.

Suppose we are given $f_{..n}$, $r_{..n}$, $\rho_{..n}$ and
$s_{..n-1}$. Then we can compute $f_{..2m}$ using
$(40.5+o(1))F(m)$ operations. This bound is contributed by the
following parts:

$(10.5+o(1))F(m)$ operations to compute $s_{..m-1}$ and DFTs of
blocks of $s_{..m-1}$ via~(7). We use DFTs of blocks of $h$, $s$,
$\Delta h$, $\rho$ and inverse DFTs to restore triple
multiplications (it is essential that the blocks of
$\Delta(h_{..m})$ are double-sized);

$(10+o(1))F(m)$ operations to compute another half of
$s_{..2m-1}$. Here each iteration~(7) is performed via two
ordinary multiplications with the use of DFTs of order $2k$;

$(10+o(1))F(m)$ operations to compute $f_{..m}$ and DTFs of its
blocks by~(5). The procedure is the same as in the first part of
the exponential algorithm;

$(6+o(1))F(m)$ operations to compute $\ln f_{..m}$ up to order
$2m$. This step coincides with that of the exponential algorithm
above;

$(4+o(1))F(m)$ operations to perform final multiplication in~(6)
via DFTs of order $2k$.

Therefore, we have got the required complexity estimate.

Research supported in part by RFBR, grants 11--01--00508,
11--01--00792, and OMN RAS ``Algebraic and combinatorial methods
of mathematical cybernetics and information systems of new
generation'' program (project \linebreak ``Problems of optimal
synthesis of control systems'').

\section*{Appendix}

Let $f,g,h \in \mathbb C[[x]]$. Consider a problem of computing
$$ q = f \lfloor g h / x^m \rfloor \mod x^n, $$
where $m$ is a multiple of $n$. Suppose series $f,g,h \in \mathbb
C[[x]]$ are given divided into blocks of size $k$ (we assume for
simplicity that $n$ is a multiple of $k$):
$$  f = \sum_{i \ge 0} a_i x^{ik}, \quad g = \sum_{i \ge 0} b_i x^{ik}, \quad h = \sum_{i \ge 0} c_i x^{ik}, \quad \deg a_i,b_i,c_i < k. $$
 Suppose we are also given DFTs of order $3k$ (or,
alternatively, double DFTs of order $(l_1,l_2)$ with $l_1+l_2=3k$)
of all necessary blocks $a_i$, $b_i$, $c_i$. We are to show how
one can compute $q$ via approximately $3(n/k)F(k)$ extra
operations.

Flooring makes some complication. We avoid it as following. Let
$$ u_i = \sum_{\mu+\nu=m/k+i} b_{\mu}c_{\nu}, \qquad \theta = \lfloor u_{-1} / x^k \rfloor. $$
Then
$$ \lfloor g h / x^m \rfloor = \theta + \sum_{i \ge 0} u_i x^{ik}. $$
Finally we have
$$ q = \sum_{i \ge 0} d_i x^{ik} \mod x^n, \qquad d_i = a_i \theta + \sum_{\lambda+\mu=i} a_{\lambda}u_{\mu}. $$
Note that $d_i$ are the polynomials of degree $<3k$.

Turn to calculations. Let $a^*$ to denote the vector of DFT (or
double DFT) of polynomial $a(x)$.

$(i)$ Compute $u^*_i$ for $i=-1,\ldots,n/k-1$. It costs
$O(k)F((m+n)/k) = O((m+n)\log((m+n)/k))$ since $u_i^*$ are the
components of the convolution of vectors $(b_0^*, b_1^*, \ldots ,
b_{(m+n)/k-1}^*)$ and $(c_0^*, c_1^*, \ldots , c_{(m+n)/k-1}^*)$.
Recall that $b_i^*$ and $c_i^*$ are the elements of the space
$\mathbb C^{3k}$ of DFT-images with component-wise multiplication.

$(ii)$ Compute $u_{-1}$ and hence $\theta$ via inverse DFT. It
costs $3F(k)+O(k)$ as far as $u_{-1}^*$ is known (in fact,
$2k$-point inverse DFT suffices here).

$(iii)$ Compute $\theta^*$. it costs $3F(k)+O(k)$.

$(iv)$ Compute $d^*_i$ for $i=0,\ldots,n/k-1$. It costs
$O(k)F(n/k) =$ \linebreak $O(n\log(n/k))$ contributed by a
convolution of order $n/k$ in $\mathbb C^{3k}$ and $O(n/k)$ extra
additions and multiplications in the same space.

$(v)$ Compute all $d_i$ and hence $q$. It costs
$(n/k)(3F(k)+O(k))$ to compute $d_i$ and $O(n)$ to restore $q$.

Finally we have an upper bound
$$3(n/k+2)F(k)+O((m+n)\log((m+n)/k))$$
for the total complexity.

\end{document}